OpenStudy (anonymous):

Simplify -(3a-2b) - 3(-a-b)

OpenStudy (anonymous):

Hint: distribute the negatives and combine like terms

OpenStudy (anonymous):

What do you mean by distribute the negatives?

OpenStudy (anonymous):

-×3a -×-2b -3×-a -3×-b

OpenStudy (anonymous):

in the first parenthesis there is a negative that can be distributed

OpenStudy (cj49):

-3a+2b+3a+3b

OpenStudy (anonymous):

-(3a-2b) ^

OpenStudy (anonymous):

behind every operation there is a number in this case its a -1 thats being distributed

OpenStudy (anonymous):

6a-1b

OpenStudy (anonymous):

i get something different -(3a-b)+3a-3b -3a+b+3a-3b now combine like terms 3a-3a -3b+1b?

OpenStudy (anonymous):

oops -3b+2b

OpenStudy (anonymous):

When I distributed I got -3a+2b+3a+3b

OpenStudy (anonymous):

Then subtracted 2 b from 3b

OpenStudy (anonymous):

and got -3a+3a-1b

OpenStudy (anonymous):

noo dont subtract when combining like terms we keep their operations if it was -3a+2b=3a+3b we would subtract but in this case it's on the same side of the equal sign so we keep the operation the same

OpenStudy (anonymous):

oh so you dont simplify anymore?

OpenStudy (anonymous):

yes you continue to simplify but keep the operations the same from here -3a+3a 2b+3b

OpenStudy (anonymous):

im confused again

OpenStudy (anonymous):

the expression is on the same side of the equal sign so we do not do their reciprocal we keep the operations the same

OpenStudy (anonymous):

http://www.mathsisfun.com/algebra/like-terms.html heres a site that might explain combining like terms better than i can

OpenStudy (anonymous):

ok i understand more now but can you show me how to combine for this problem?

OpenStudy (anonymous):

-a+a=0 b+b=2b heres an example

OpenStudy (anonymous):

i was just confused about the 0

OpenStudy (anonymous):

theres no exponent for 0 is there? its just 0 right

OpenStudy (anonymous):

yes it's just 0 you dont have to write a term for it

OpenStudy (anonymous):

so it would be 0 + 5b^22 ??

OpenStudy (anonymous):

i mean 2

OpenStudy (anonymous):

not 22

OpenStudy (anonymous):

the only time you would write a exponent with the variable is if you multiply the same variables together in this case we dont have an exponent so 5b would be the answer

OpenStudy (anonymous):

ok!! thank you so much!

OpenStudy (anonymous):

you're welcome